Subject: APO 3.5m users committee minutes, 2/8/10

From: Suzanne Hawley

Submitted: Tue, 9 Feb 2010 13:31:26 -0800 (PST)

Message number: 1168 (previous: 1167, next: 1169 up: Index)

          APO 3.5-m Users Committee Phonecon, 2/8/10

Attending: Suzanne Hawley, Mark Klaene, Bill Ketzeback, Al Harper, 
John Bally, Michael Strauss, Jon Holtzman, Scott Anderson, 
Jon Fulbright

Absent:  Remy Indebetouw


User feedback, comments from institutional representatives:

Chicago (Al Harper) - nothing to report.

New Mexico State (Jon Holtzman) - Would be useful to know which TUI 
software version updates are critical and which are simply desired.
Latest update had issues with installation on linux machine.

Washington (Scott Anderson) - nothing to report.

Colorado (John Bally) - nothing to report.

Princeton (Michael Strauss) - nothing to report.

Johns Hopkins (Jon Fulbright) - nothing to report.

Virginia - no report (Indebetouw absent)


Discussion of telescope/instruments report - Klaene

Mark called attention to some recent problems associated
with the telescope, including magna-sensor errors and altitude 
drive preload. Both of these appear to be relatively minor and 
have cost little time. Major loss time factor has been the weather.


3.5-m Telescope, Instruments, and CIF Projects Highlights, 1/4/10 through 2/2/10

0) Overview

Weather, specifically snow, has been the main issue again this period 
and we expect still more.  Large snow falls have created a lot of staff 
access problems as well as power losses.  Loss time due to snow on the 
roof was minimal, however high and variable humidity continues to keep 
us closed on some occasions under clear skies.

1) Telescope

Behavior has been quite good with only rare minor issues with pointing 
and tracking. We have also seen very few cold-temperature related problems.

2) Instruments

DIS, Echelle, SPIcam, Agile and TripleSpec ran with no problems. 
DIS blue ION pump is still running a little high.

The DIS slitviewer camera shutter had its usual cold weather problems 
last period. We are monitoring this for increasing frequency but so far 
it has not re-occurred.

The new NICFPS fanout board is progressing slowly with the board at CU 
for final assembly and testing.  Installation is now set for March. 
There was still no progress on the NICFPS etalon at CU.

3) CIF projects

High altitude tracking data are being analyzed. The spare drive box 
design is complete and fabrication has started.


o NICFPS fanout board, possible scheduling impact - Klaene

After some discussion it was decided to test the new fanout board with 
the engineering chip at CU before installation at APO. This will delay 
any installation until March at the earliest.  Once again the plan is 
to perform the installation during time when NICFPS is not scheduled to
be on the telescope. John Bally knows of no progress on the F-P Etalon. 
Could be tested this spring.

o Triplespec scheduling in Q2 - Hawley

Feedback from the Observing Specialists is that they are getting more
familiar with mounting Triplespec even during the cold winter months,
and therefore block scheduling restrictions will be relaxed for the
remainder of Q1 and into Q2.  However users should be aware that 
Triplespec instrument changes do take longer than the other 
instruments which directly translates to a loss of observing time 
on the sky if requests are made during observing (ie changing instrument
in the middle of a half night).

o Q2 scheduling and class trips - Hawley

CU class scheduled for March, UW class for July.  Currently we have none 
scheduled for Q2 but if there are any plans for class trips to APO in Q2 
they need to coordinate scheduling with Mark and Suzanne. Housing on site 
is still limited and needs to be worked out well in advance.  Some progress
has been made on obtaining new single beds to augment space in the dorm

o Echelle update - Hawley, Holtzman

Jon Holtzman has agreed to lead the echelle upgrade project, with help 
from JHU and possibly UChicago personnel. Work will begin on a conceptual 
design and requirements document with the aim of having a PDR in fall 2010.



[open from previous months]: NONE

[new actions from this meeting]: NONE


Next meeting: The next Users Committee phonecon will be on Monday, 8 March, 
at 8:30 AM Pacific Time.  The agenda and other materials will be sent 
to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1168 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to